OpenStudy (anonymous):

solve ysec2x y'=-x(1-y^2)

OpenStudy (abb0t):

I'm going to assume this is an IVP? Correct?

OpenStudy (anonymous):

\[y \sec2x \frac{ dy }{ dx}=-x \sqrt{1-y ^{2})}\]

OpenStudy (abb0t):

Separate the x's and y's on each side and take the integral of each

OpenStudy (abb0t):

\[\int\limits \frac{y }{ 1+y^2} dy = \int\limits \frac{ x }{ \sec(2x)}dx\]

OpenStudy (abb0t):

If you aren't given any initial conditions, then you can leave your answer in the form f(y) = g(x) + c

OpenStudy (anonymous):

why did the 1-y^2 turn to 1+y^2

OpenStudy (abb0t):

Typo. My bad homie It should be: 1-y^2

OpenStudy (anonymous):

i am coming up with \[y=\sqrt{1-(\frac{ x }{2}}\sin 2x +1/4\cos2x+c)^{2}\]

OpenStudy (anonymous):

would you think that is realistic?

OpenStudy (abb0t):

no.

OpenStudy (abb0t):

Is this the problem you are solving for: \[y \sec(2x)\frac{ dy }{ dx } = x (1-y^2)\] correct?

OpenStudy (abb0t):

What you have there is a nonlinear first order differential equation. and since you aren't given any initial conditions, you can leave your answer in implicit form.

OpenStudy (anonymous):

i am given \[y(\pi)=1\]

OpenStudy (abb0t):

That means that y = 1 and x = π So just substitute and solve for 'C' once you have your C, plug it back into your general solution to get your explicit solution. Did you figure out your general solution?

OpenStudy (anonymous):

what i typed ealier is what i am getting for y

OpenStudy (anonymous):

i am getting c as -1/4

OpenStudy (abb0t):

I don't see how you got that. Your implicit solution should be: \[\frac{ 1 }{ 2 }\ln(y^2-1) = -[ xsin(x)+\cos(x)+c] \]

OpenStudy (anonymous):

when i separate them this is what i got\[\int\limits_{}^{}\frac{ y }{\sqrt{(1-y^2}}dy=\int\limits_{}^{}xcos2xdx\]

OpenStudy (abb0t):

I didn't know there was a square root. My answer is wrong then Let me re-work it. Give me a minute.

OpenStudy (abb0t):

but that looks correct. the way you separated them!

OpenStudy (abb0t):

Solution: \[-\sqrt{1-y^2}= \frac{ 1 }{ 4 }(-2xsin(2x)-\cos(2x))\]

OpenStudy (abb0t):

C = -1

OpenStudy (anonymous):

we can multiply by -1 through out? and you forgot the c

OpenStudy (abb0t):

Yeah. I 4got to add +C. Good catch! So your solution should be: \[-\sqrt{1-y^2} = \frac{ 1 }{ 4 }[-2xsin(2x)-\cos(2x)-1]\]

OpenStudy (anonymous):

this is what i got before i further simplified

OpenStudy (anonymous):

\[\sqrt{1-y ^{2}}=\frac{ x }{ 2 }\sin2x +\frac{ 1 }{ 4 }\cos2x+c\]

OpenStudy (abb0t):

That's correct.

OpenStudy (anonymous):

sub x=pi and y=1 RHS=0 sin2x term becomes zero leaving 1/4+c=0

OpenStudy (abb0t):

You're correct. I 4got abt the 1/4. My bad. Haha. But you get it better than me, since u caught my mistakes :)

OpenStudy (anonymous):

no, it took a while to do it and u was doin it as we were talkin so u r better

OpenStudy (anonymous):

and u juss confirmed my answers. thx

OpenStudy (abb0t):

no prb. best of luck.